OpenStudy (anonymous):

Are these functions inverses of each other? f(x) = x + 6 f(x) = x - 6

OpenStudy (anonymous):

If you plug one into the other you should just get x.

OpenStudy (anonymous):

If it works both ways, then they are inverses.

OpenStudy (anonymous):

so no

OpenStudy (anonymous):

Really? f(x-6) = (x-6) + 6 = x - 6 + 6 = x

OpenStudy (anonymous):

So that one is x

OpenStudy (anonymous):

for the other one.. \[f_2(x+6) = (x+6) - 6 = x + 6 - 6 = x\]

OpenStudy (anonymous):

So yeah, they both equal x when you plug in the other one for x. Therefore they are inverses.

OpenStudy (anonymous):

thank you so much!